# Constrained Portfolio Analysis in High Dimensions: Tracking Error and Weight Constraints

### v1

@@ 7 Empirical study
We report the average returns (AVR), tracking error (TE), standard deviation (Risk), and Sharpe ratio (SR). Following Ao et al. (2019) and Caner et al. (2023), we conduct hypothesis tests regarding the Sharpe ratio to check the statistical difference of different models. Specifically, we test
$H_{0}\mathrel{\mathop{:}}SR_{rb-NW}\leqslant SR_{0}\quad\text{ vs }\quad H_{a}\mathrel{\mathop{:}}SR_{rb-NW}>SR_{0}$ (12)
where $SR_{rb-NW}$ denotes the Sharpe ratio of the rb-NW portfolio, and $SR_{0}$ denotes the Sharpe ratio of the portfolio under comparison such as POET, SFNL, etc. We report the p-values of the Sharpe ratio test (Memmel, 2003), which corrects the test of Jobson and Korkie (1981). We also used the Ledoit and Wolf (2008) test with circular bootstrap, and the results were very similar; therefore we only report those of the first test in the following tables.
We present the results based on monthly and annual returns (those are reported in the subpanels of “Monthly” and “Yearly” of Table 7 and the following empirical tables), respectively, following the practice of the fund report. By yearly we take averages of 30 annual returns in out-sample, whereas in monthly we take averages of 360 monthly returns in out-sample.
In scenario (1), we find rb-NW achieves the best return/risk balance. Using the monthly result in Table 7 as an example, the Sharpe ratio of rb-NW portfolio is 0.2542, with an average return of 0.0103 and risk of 0.0407, all top performers among the comparison strategies. The results of scenario (2) are similar, such that the rb-NW portfolio obtains the highest average return and Sharpe ratio. It has the best performance in all measures using yearly returns. The p-values of the comparison test in (12) suggest that the proposed rb-NW portfolio generally has statistically significantly better result performance among comparison portfolios when $p<T$. For $p>T$ case, the p-values show that the SR performance of rb-NW is not statistically significant from other methods. Figures 1 and 2 visualize the cumulative returns of comparison portfolios for scenarios (1) and (2), respectively. Similar to what Table 7 has shown, it can be observed that the curve of rb-NW based portfolio outperforms other portfolios in the studied period. NW-based portfolio and SFNL-based portfolio rank second in scenarios (1) and (2), respectively. The difference between rb-NW based portfolio and the benchmark market index is economically significant. To substantiate that in Figure 1, our method achieves almost 3000% cumulative return between 1991-January to December 2020, whereas index stays at 1000% cumulative return.
Table 7: Empirical results for scenarios (1) and (2).
Scenario (1): p = 100, T = 120 Scenario (2): p = 254, T =180
AVR TE Risk SR p AVR TE Risk SR p
Monthly
Index 0.0077 - 0.0420 0.1826 0.0000 0.0071 - 0.0437 0.1627 0.0067
NW 0.0101 0.0132 0.0423 0.2381 0.0205 0.0087 0.0241 0.0440 0.1970 0.0259
rb-NW 0.0103 0.0103 0.0407 0.2542 - 0.0093 0.0188 0.0406 0.2296 -
POET 0.0100 0.0189 0.0419 0.2379 0.0995 0.0094 0.0353 0.0434 0.2178 0.2525
NLS 0.0097 0.0193 0.0426 0.2280 0.0313 0.0092 0.0274 0.0421 0.2189 0.3221
SFNL 0.0101 0.0146 0.0420 0.2396 0.0586 0.0093 0.0244 0.0418 0.2218 0.3392
Yearly
Index 0.0965 - 0.1713 0.5634 0.0086 0.0858 - 0.1729 0.4963 0.0200
NW 0.1263 0.0500 0.1636 0.7720 0.1153 0.0983 0.0992 0.1633 0.6020 0.0833
rb-NW 0.1294 0.0384 0.1607 0.8054 - 0.1085 0.0723 0.1550 0.6996 -
POET 0.1204 0.0750 0.1693 0.7114 0.0932 0.1004 0.1609 0.1666 0.6025 0.3082
NLS 0.1174 0.0809 0.1718 0.6836 0.0605 0.1019 0.1238 0.1598 0.6375 0.3132
SFNL 0.1239 0.0603 0.1627 0.7612 0.1763 0.1039 0.1147 0.1574 0.6603 0.3580
  • We report the following statistics for portfolio performance evaluations: average returns (AVR), tracking error (TE), standard deviation (Risk) and Sharpe ratio (SR). “p” is the p-value for the SR test. “Index” is the S&P 500 market index, for both scenarios. Other portfolios are explained in Table 1 notes.
Figure 1: Cumulative excess return of largest 100 stocks in S&P 500 index, out of sample period is 01/1991-12/2020, sample size is 120.
Figure 2: Cumulative excess return of 254 stocks in S&P 500 index, out of sample period is 01/2006-12/2020, sample size is 180.
We also take transaction costs into consideration. Following Ao et al. (2019), the excess portfolio return net of transaction cost is computed as
$r_{t}^{net}=\left(1-\sum_{i=1}^{p}c\left|\widehat{W}_{t+1,i}-\widehat{W}_{t,i}^{+}\right|\right)(1+r_{t})-1,$ (13)
where $\widehat{W}_{t+1,i}$ is the $i$-th element of portfolio weight after rebalancing, and $\widehat{W}_{t,i}^{+}$ is the portfolio weight before rebalancing; $r_{t}$ is the excess return of the portfolio without transaction cost and $c$ measures the level of the transaction cost, which is set to 50 basis points. Formally, the portfolio turnover is defined as
$TO=\frac{1}{RT}\sum_{l=1}^{RT}\sum_{i=1}^{p}\left|\widehat{W}_{l+1,i}-\widehat{W}_{l,i}^{+}\right|,$ (14)
where $\widehat{W}_{l+1,i}$ is the desired portfolio weight at $(l+1)$-th rebalancing, and $\widehat{W}_{l,i}^{+}$ is portfolio weight before the $(l+1)$-th rebalancing, $RT$ is the number of rebalancing events. The empirical results for scenarios (1) and (2) are reported in Table 8. It shows that the advantages of our proposed portfolio achieved without transaction costs still hold with transaction costs. E.g., the Sharpe ratio of rb-NW portfolio is a bit lower with transaction costs compared to Table 7, but it is still the best among comparison portfolios, and the advantages are more significant when $p>T$. Notice that rb-NW based portfolio gets the lowest turnover ratio, as it has the lowest tracking error to the benchmark index. The p-values of SR test results show that the rb-NW has better SR performance compared to other portfolios when $p<T$, and similar outcome is observed for $p>T$ and monthly performance, at 10% for example.
Table 8: Empirical results for scenarios (1) and (2) with transaction cost.
Scenario (1): p = 100, T = 120 Scenario (2): p = 254, T = 180
AVR TE Risk SR TO p AVR TE Risk SR TO p
Monthly
Index 0.0077 - 0.0420 0.1826 - - 0.0071 - 0.0437 0.1627 -
NW 0.0093 0.0132 0.0423 0.2189 0.1638 0.0032 0.0069 0.0242 0.0439 0.1578 0.3455 0.0029
rb-NW 0.0098 0.0103 0.0407 0.2404 0.1125 - 0.0083 0.0188 0.0406 0.2046 0.2045 -
POET 0.0086 0.0190 0.0419 0.2047 0.2784 0.0026 0.0070 0.0353 0.0433 0.1606 0.4971 0.0862
NLS 0.0082 0.0195 0.0427 0.1932 0.2951 0.0000 0.0072 0.0277 0.0421 0.1704 0.4070 0.0745
SFNL 0.0089 0.0147 0.0420 0.2126 0.2256 0.0017 0.0075 0.0245 0.0418 0.1788 0.3599 0.0894
Yearly
Index 0.0965 - 0.1713 0.5634 - - 0.0858 - 0.1729 0.4963 - -
NW 0.1154 0.0493 0.1617 0.7135 1.9817 0.0380 0.0762 0.0969 0.1596 0.4772 4.2072 0.0220
rb-NW 0.1220 0.0377 0.1595 0.7650 1.3600 - 0.0953 0.0713 0.1528 0.6236 2.4660 -
POET 0.1022 0.0733 0.1657 0.6169 3.3184 0.0170 0.0686 0.1556 0.1589 0.4315 5.9400 0.1610
NLS 0.0980 0.0800 0.1684 0.5817 3.5734 0.0112 0.0755 0.1222 0.1553 0.4860 5.0076 0.1415
SFNL 0.1088 0.0600 0.1603 0.6791 2.7319 0.0345 0.0806 0.1128 0.1526 0.5280 4.3788 0.1807
  • TO is the turnover defined in (14). For other measures please see notes in Table 7.
Next, we conduct empirical exercises to study the portfolios with both TE and inequality weight constraints, mimicking the mutual funds focusing on specific sectors (themes). Here, we consider two sectors, the health care sector and the energy sector, and for ease of reporting, we call them scenarios (3) and (4), respectively. Specifically, we select the health care/energy assets from the same S&P 500 assets pool in scenario (2) with 254 stocks, and impose the weight constraint such that our portfolio invests at least 80% of its assets in the themed assets. To compare with TE-only constraint of scenario (2), the out of sample period is set to 01/2006-12/2020 with sample size $T=180$. The benchmark indices for scenarios (3) and (4) are S&P 500 Health Care and S&P 500 Energy, respectively. The results are shown in Table 9. We can observe that the rb-NW based portfolios show overall outstanding performances in average return, risk, and Sharpe ratio. First, the Sharpe ratio of the proposed rb-NW is the best among all portfolios in scenarios (3) and (4). For example, under the ”Yearly” scenario (3), the Sharpe ratio of rb-NW based portfolio is 0.8853, which is higher than the second-ranked POET (0.8227). The shrinkage-based portfolios have outperformed the respective sector indices. We can also compare the themed funds with the market index: from scenario (2) of Table 7, which has the same stocks pool and time period, the market index of S&P 500 has better Sharpe ratio than the energy funds, while the health care funds have higher Sharpe ratio than the market index. Second, the risk of rb-NW portfolios in scenario (3) is the best among all active portfolios and is close to the best (Index) in energy funds. At last, the average returns of rb-NW based portfolios are also the top performer in the two themed funds. We also illustrate the superior returns of rb-NW in Figures 3 and 4, which is consistent with the average returns result in Table 9.
Table 9: Empirical results for scenarios (3) and (4).
Scenario (3): p = 254, T =180, Health Care Scenario (4): p = 254, T =180, Energy
AVR TE Risk SR p AVR TE Risk SR p
Monthly
Index 0.0079 - 0.0400 0.1988 0.1308 0.0012 - 0.0724 0.0168 0.0165
NW 0.0084 0.0239 0.0414 0.2036 0.0430 0.0045 0.0361 0.0825 0.0546 0.0974
rb-NW 0.0093 0.0128 0.0383 0.2435 - 0.0056 0.0224 0.0768 0.0733 -
POET 0.0091 0.0179 0.0385 0.2376 0.3418 0.0044 0.0494 0.0756 0.0580 0.3873
NLS 0.0084 0.0193 0.0401 0.2096 0.0257 0.0037 0.0362 0.0780 0.0480 0.0873
SFNL 0.0088 0.0172 0.0398 0.2223 0.0885 0.0049 0.0285 0.0778 0.0628 0.3742
Yearly
Index 0.0973 - 0.1488 0.6543 0.0659 0.0349 - 0.2034 0.1716 0.0433
NW 0.0946 0.0973 0.1353 0.6991 0.0826 0.0360 0.1521 0.2214 0.1628 0.0966
rb-NW 0.1086 0.0501 0.1227 0.8853 - 0.0561 0.0924 0.2020 0.2778 -
POET 0.1049 0.0944 0.1275 0.8227 0.3224 0.0340 0.2534 0.2190 0.1552 0.4280
NLS 0.0925 0.1004 0.1353 0.6834 0.0931 0.0223 0.1934 0.1973 0.1130 0.2268
SFNL 0.0983 0.0969 0.1306 0.7526 0.1807 0.0383 0.1652 0.1858 0.2064 0.4397
  • We report the following statistics for portfolio performance evaluations: average returns (AVR), tracking error (TE), standard deviation (Risk), Sharpe ratio (SR) and p-value of the SR test. Scenarios (3) and (4) are themed portfolios (composed of 254 stocks from 01/2006 - 12/2020 from S&P500 index) benchmarking against health care and energy sector, respectively. “-” means the statistic is not available. “Index” means the S&P 500 Health Care and S&P 500 Energy indices, for scenario (3) and (4), respectively. Other portfolios are explained in Table 1 notes.
Figure 3: Cumulative excess return of all considered portfolios under scenario (3) (health care sector), out of sample period is 01/2006-12/2020, sample size is 180.
Figure 4: Cumulative excess return of all considered portfolios under scenario (4) (energy sector), out of sample period is 01/2006-12/2020, sample size is 180.
The empirical results for scenarios (3) and (4) with transaction costs are reported in Table 10. It shows that the robustness of the superior performance of the proposed portfolio with transaction costs in the two designated sectors. Same as Table 9, the rb-NW portfolio is still the best performer in health care sector after transaction costs. It has good performance in the energy sector in terms of Sharpe ratio. The results of the SR tests support the better performance of rb-NW in general.
Table 10: Empirical results for scenarios (3) and (4) with transaction cost.
Scenario (3): p = 254, T =180, Health Care Scenario (4): p = 254, T =180, Energy
AVR TE Risk SR TO p AVR TE Risk SR TO p
Monthly
Index 0.0079 - 0.0400 0.1988 - - 0.0012 - 0.0724 0.0168 - -
NW 0.0069 0.0241 0.0415 0.1668 0.3038 0.0083 0.0031 0.0291 0.0810 0.0380 0.3929 0.0282
rb-NW 0.0085 0.0129 0.0383 0.2229 0.1570 - 0.0042 0.0225 0.0767 0.0547 0.2870 -
POET 0.0071 0.0180 0.0385 0.1856 0.3983 0.0057 0.0009 0.0373 0.0742 0.0119 0.8448 0.0081
NLS 0.0067 0.0196 0.0401 0.1678 0.3324 0.0011 0.0020 0.0283 0.0772 0.0260 0.5051 0.0056
SFNL 0.0073 0.0174 0.0398 0.1834 0.3081 0.0068 0.0035 0.0213 0.0776 0.0449 0.3977 0.1471
Yearly
Index 0.0973 - 0.1488 0.6543 - - 0.0349 - 0.2034 0.1716 - -
NW 0.0752 0.0975 0.1342 0.5606 3.7668 0.0335 0.0220 0.1237 0.2124 0.1035 4.8096 0.0402
rb-NW 0.0985 0.0507 0.1218 0.8086 1.9368 - 0.0392 0.0938 0.2008 0.1952 3.4476 -
POET 0.0803 0.0938 0.1233 0.6513 4.6104 0.1082 -0.0033 0.1939 0.1930 -0.0171 9.6120 0.0829
NLS 0.0713 0.1013 0.1338 0.5329 4.1316 0.0393 0.0053 0.1545 0.1865 0.0282 6.1596 0.0703
SFNL 0.0785 0.0976 0.1285 0.6108 3.8076 0.0894 0.0254 0.1270 0.1828 0.1387 4.7988 0.1654
@@
8 Conclusion
In this paper we develop the first high dimensional constrained portfolio weight estimator by using a new machine learning technique. We prove that our estimator is consistent even when the number of assets are larger than the time span of the portfolio and when the number of factors may grow. We provide rate of convergence results when there are tracking error constraints, or joint tracking error and inequality-equality weight constraints. The future research may look at confidence intervals for Sharpe Ratio of the constrained portfolio.

### v2

@@ 7 Empirical Study
Following Ao et al. (2019) and Caner et al. (2023), we conduct hypothesis tests regarding the Sharpe Ratio to check the statistical difference between different models. For the Sharpe Ratio test, specifically, we test
$H_{0}:SR_{CROWN}\leq SR_{0}\quad\text{ vs }\quad H_{a}:SR_{CROWN}>SR_{0}$ (12)
where $SR_{CROWN}$ denotes the Sharpe Ratio of the CROWN portfolio, and $SR_{0}$ denotes the Sharpe Ratio of the portfolio under comparison such as POET, SFNL. We report the p-values of the Sharpe Ratio test (Ledoit and Wolf, 2008), and the results were very similar to the test of Memmel (2003). We also consider transaction costs. Following Ao et al. (2019), the excess portfolio return net of transaction cost is computed as $y_{t}^{net}=\left(1-\sum_{i=1}^{p}c\left|\widehat{w}_{t+1,i}-\widehat{w}_{t,i}^{+}\right|\right)(1+y_{t})-1,$ where $\widehat{w}_{t+1,i}$ is the $i$ th element of portfolio weight after rebalancing, and $\widehat{w}_{t,i}^{+}$ is the portfolio weight before rebalancing; $\boldsymbol{y}_{t}$ is the excess return of the portfolio without transaction cost and $c$ measures the level of the transaction cost, which is set to 50 basis points. Formally, the portfolio turnover is defined as $TO=\frac{1}{RT}\sum_{l=1}^{RT}\sum_{i=1}^{p}\left|\widehat{w}_{l+1,i}-\widehat{w}_{l,i}^{+}\right|,$ where $\widehat{w}_{l+1,i}$ is the desired portfolio weight at $(l+1)$ th rebalancing, and $\widehat{w}_{l,i}^{+}$ is portfolio weight before the $(l+1)$ th rebalancing, $RT$ is the number of rebalancing events. Note that all our Tables 5-7 have taken into account transaction costs. At last, the targeted TE is 5% annualized.
To evaluate the performance of those implementable portfolios, we compute the infeasible out-of-sample oracle portfolio. For this oracle portfolio, instead of returns in the backward-looking sample period, the expected return $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}_{y}$ are estimated with returns of the stocks in the out-of-sample periods (01/1996-12/2020 for S&P 500 Index and 01/2006-12/2020 for S&P 500 Sector indices). Our comparison group includes the benchmark, INDEX, which is the market index portfolio, and the infeasible out-of-sample oracle portfolio.We also compare CROWN with other popular method for covariance/precision matrix estimation detailed in the simulation section. In the TE constraint case, CROWN achieves the best return/risk balance. As shown in Table 5, the Sharpe Ratio of CROWN portfolio is 0.0938, whereas POET has 0.0828, and NLS has 0.0676. Notice that our proposed portfolio has the lowest turnover ratio, as it has the lowest TE to the benchmark index. The p-values of the SR comparison test suggest that the proposed CROWN portfolio has statistically significantly better performance at 1% level among comparison portfolios, after considering the transaction cost. The result shows that as $p>T_{I}$, in Table 5, the proposed CROWN performs well as suggested by the theory.
Table 5: Empirical results for TE constraint
S&P 500 Index
AVR TE Risk SR TO p-val Max Min Total
Weight Weight Short
Oracle 0.0180 0.0109 0.0432 0.4162 0.1072 0.0000 0.0512 -0.0002 0.0012
INDEX 0.0037 - 0.0418 0.0894 - 0.0000 - - -
NW 0.0034 0.0037 0.0417 0.0827 0.6960 0.0038 0.0944 -0.0452 1.0634
CROWN 0.0039 0.0021 0.0418 0.0938 0.6096 - 0.0417 -0.0000 0.0000
POET 0.0035 0.0032 0.0417 0.0828 0.6934 0.0000 4.6343 -4.0492 160.0132
NLS 0.0028 0.0046 0.0412 0.0676 0.7989 0.0000 0.1482 -0.0385 1.6725
SFNL 0.0033 0.0033 0.0414 0.0799 0.7129 0.0000 0.0369 -0.0020 1.6129
  • Notations: average returns (AVR), tracking error (TE), standard deviation (Risk), Sharpe Ratio (SR), turnover ratio (TO). “p-val” is the p-value for the SR test. “Max/Min weight” is the maximum/minimum individual asset weight in the portfolio, “Total short” is the total shorting position of the portfolio. “Oracle” is the infeasible out-of-sample oracle portfolio, “Index” refers to S&P 500 Index, S&P 500 Healthcare Sector Index, and S&P 500 Energy Sector Index in Table 5, Table 6, and Table 7, respectively.
Table 6: Empirical results with TE and inequality weight constraint
S&P Healthcare Index
AVR TE Risk SR TO p-val %Themed Max Min Total
Assets Weight Weight Short
Oracle 0.0120 0.0039 0.0385 0.3118 0.0408 0.0000 0.8000 0.1297 -0.0002 0.0020
INDEX 0.0081 - 0.0389 0.2088 - 0.4928 0.8000 - - -
NW 0.0080 0.0023 0.0384 0.2076 0.0673 0.4244 0.8031 0.1201 -0.0019 0.0474
CROWN 0.0078 0.0059 0.0373 0.2090 0.0853 - 0.8035 0.1201 -0.0092 0.1097
POET 0.0058 0.0057 0.0379 0.1541 0.5344 0.0000 0.8019 0.1212 -0.0042 0.1105
NLS 0.0077 0.0033 0.0382 0.2027 0.0879 0.1689 0.8036 0.1211 -0.0052 0.1226
SFNL 0.0079 0.0019 0.0384 0.2056 0.0703 0.3452 0.8017 0.1208 -0.0030 0.0650
@@
Table 7: Empirical results with TE and inequality weight constraint
S&P 500 Energy Index
AVR TE Risk SR TO p-val %Themed Max Min Total
Assets Weight Weight Short
Oracle 0.0090 0.0167 0.0597 0.1504 0.0364 0.0000 0.8000 0.2945 -0.0029 0.1261
INDEX 0.0034 - 0.0723 0.0473 - 0.1586 0.8000 - - -
NW 0.0031 0.0157 0.0601 0.0513 0.1131 0.2337 0.8045 0.2868 -0.0055 0.1903
CROWN 0.0035 0.0135 0.0619 0.0569 0.0396 - 0.8001 0.2883 -0.0000 0.0002
POET 0.0031 0.0134 0.0624 0.0497 0.1327 0.0018 0.8001 0.2884 -0.0006 0.0173
NLS 0.0031 0.0147 0.0605 0.0508 0.0906 0.0711 0.8022 0.2877 -0.0036 0.1477
SFNL 0.0033 0.0140 0.0612 0.0534 0.0704 0.0723 0.8016 0.2880 -0.0021 0.0793
  • See notes in Table 5.
Next, we conduct empirical exercises to study the portfolios with both TE and inequality weight constraints. Specifically, we select the healthcare/energy assets from the same S&P 500 asset pool, and impose the weight constraint such that our portfolio invests at least 80% of its assets in the themed assets. The number of energy sector stocks is 24, and the number of healthcare stocks is 49. The results are shown in Tables 6 and 7. We can observe that the proposed portfolios (CROWN) show overall good performances in average return, risk, and Sharpe Ratio. First, the Sharpe Ratio of the proposed CROWN is the best among all portfolios in both cases. In the healthcare case, our method also has the lowest risk as well among all methods. For the healthcare index case, the SR tests generally do not reject the null of indifferent performance. For the energy index case, the p-values are mostly significant at the 10% level, suggesting the statistical significance of the CROWN’s better performance. The shrinkage-based portfolios also outperformed the energy index. Finally, the average returns of our proposed portfolio is also the top performer in the energy sector.
8 Conclusion and Discussion
In this paper, we pioneer the development of the high-dimensional constrained portfolio weights estimator using a novel statistical technique. Our analysis demonstrates the consistency of the estimator, even in scenarios where the number of assets surpasses the portfolio’s time span. We provide rate of convergence results for scenarios involving TE constraints or joint TE and inequality-equality weight constraints.
Disclosure Statement
The authors report there are no competing interests to declare.